OpenStudy (anonymous):

URGENT The area of a rectangle is at most 21 square inches. The width of the rectangle is 3.5 inches. What are the possible measurements for the length of the rectangle?

OpenStudy (anonymous):

0<L<or equal to21/3.5

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

0<L<or equal to 6

OpenStudy (anonymous):

whats O and L ?

OpenStudy (anonymous):

0 is zero l is ur length ur length has to be greater than zero but less than or equal to 6

OpenStudy (anonymous):

because area is length*width A=L*W 21=L*(3.5) at most

OpenStudy (anonymous):

so the answer is?

OpenStudy (anonymous):

greater than 0 inches but less than or equal to 6

OpenStudy (anonymous):

\[0<L \le\]

OpenStudy (anonymous):

6

OpenStudy (anonymous):

exactly